OpenStudy (anonymous):

Which of the following polynomials does not have real roots? A. 2x2 - 13 B. x2 - 17x - 7 C. x4 + 2x2 + 2 D. -x4 + 4x2 + 6

OpenStudy (cruffo):

Have you tried using the discriminant: \(\large b^2 - 4ac\)

OpenStudy (anonymous):

C.\[\left\{x=\pm \sqrt{-1-i},x=\pm \sqrt{-1+i}\right\} \]

OpenStudy (cruffo):

Yah, that looks good.